OpenStudy (anonymous):

8x3-27 How would you factor it?

OpenStudy (abb0t):

Take the cube of both sides: \[\sqrt[3]{8x^3} = \sqrt[3]{27}\]

OpenStudy (anonymous):

Sorry i entered the question wrong! its 8x3-27 @abb0t

OpenStudy (mertsj):

Rewrite it like this: \[(2x)^3-(3)^3\] And use this pattern: \[a^3-b^3=(a-b)(a^2+ab+b^2)\]